OpenStudy (anonymous):

Rewrite this equation in standard form: `3/(x - 5) + 2 = 1/(x - 5)`

OpenStudy (phi):

the first step is multiply both sides of the equation (and *all* terms) by (x-5)

OpenStudy (phi):

the first step should give you 3 + 2(x-5) = 1 now distribute the 2, 3 + 2x -10 = 1 simplify to get 2x -7 = 1 2x -8= 0 that is "standard form" for a polynomial